The average cholesterol level for people in the United States is 210. According to the formula, how many deaths per year from heart disease can be expected with this cholesterol level? The formula given is: d + 525,000 = 5000c.
I am having trouble figuring this question out. Thanks

Answer by Adam(64)

It is a linear equation setting up a certain relationship between the level of cholesterol(c) and number of deaths per yaear from heart disease.(d)
We can manipulate the equation any way we please, but we always have to manipulate by both sides. We can subtract, add, multply or divide both sides of equation by some number.
d + 525,000 = 5000c /right from start we subtract 525,000 from both sides of equation(our goal is to seperate the variabele d.
thus
d +525,000-525,000 = 5000c-525,000
d = 5000c-525,000
We are told that average level of cholesterol is 210, so we substitute it.
d = 5000*250-525000
d = 1,250,000 - 525,000
d = 725,000
So we can expect about 725,000 deaths from heart disease per year.

By the way chloesterol is measured in milligrammes per decilitre (mg/dL)

Here are some information :
Total Cholesterol Level Category

Less than 200 mg/dL
Desirable level that puts you at lower risk for coronary heart disease. A cholesterol level of 200 mg/dL or higher raises your risk.

200 to 239 mg/dL Borderline high

240 mg/dL and above
High blood cholesterol. A person with this level has more than twice the risk of coronary heart disease as someone whose cholesterol is below 200 mg/dL.
Theese numbers look terrifing, although if you have 201 mg/dL of cholesterol in your blood probability, that you die within a 10 years from a heart disease is about 11/1200.(assuming given equation). So good luck both with cholesterol and math. :-)